OpenStudy (anonymous):

An object's speed is 2.40 m/s and its momentum is 120.0 kg * m/s. What is the mass of the object? A. 290 kg B. 0.0200 kg C. 12.0 kg D. 50.0 kg **not sure!! thank you!!

OpenStudy (michele_laino):

here we have to apply the definition of momentum P of a particle, whose mass is m, and whose velocity is v: \[P = mv\]

OpenStudy (anonymous):

ohh okay!! so 2.4 * 120? = 288? so about 290? choice A is our solution?

OpenStudy (michele_laino):

no, it is: \[\Large m = \frac{P}{v} = \frac{{120}}{{2.4}}\]

OpenStudy (anonymous):

ohh okay so 50? our solution is choice D?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

yay! thank you!!

OpenStudy (michele_laino):

thanks!! :)